OpenStudy (anonymous):

Write an equation of the ellispe with a vertex at (9,0), a co-vertex at (0,5), and center at (0,0)

OpenStudy (mertsj):

If the vertex is (9,0) then a = 9. If the co-vertex is (0,5) then b = 5 Plug into this: \[\frac{x^2}{a^2}+\frac{y^2}{b^2}=1\]

OpenStudy (aravindg):

Hi, I can see that you are a new member here at OpenStudy so, I would first like to say Welcome to OpenStudy, I would like to point you to the chat pods these chat pods are where you can make new friends and talk to new people just like you, I would also like to emphasize our "NON-CHEATING POLICY" we ask here at OpenStudy that you do NOT post exam/Test questions if you are caught doing this we will notify your current school, I would also like to ask you to check out the OpenStudy "Code of Conduct" http://openstudy.com/code-of-conduct Please read this carefully and thoroughly. Welcome to the OpenStudy Community! Ambassador

OpenStudy (anonymous):

Thanks